Is 14 Rational, Irrational, or Real?
Answer by oberobic(2304)

Rational numbers can be depicted as a fraction of integers (whole numbers), a/b provided b is not = 0.
(Dividing by 0 is undefined.)
Note that it okay for 'b' to = 1.
.
14 = 14/1 = 28/2 = 140/10 ...
.
All integers are rational numbers.
.
Done.